OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, -1) and has a slope of 5.

OpenStudy (anonymous):

where are the equations?

OpenStudy (anonymous):

y - 1 = 5(x - 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y - 2 = 5(x - 1)

OpenStudy (anonymous):

fill in the point in equations and see which equation is true

OpenStudy (anonymous):

can you help me I don't get this at all

OpenStudy (anonymous):

lets try the first one: y-1=5(x-2) fill in the point like this: -2-1=5(-2-2)

OpenStudy (anonymous):

then solve each side by itself: -3=5(-4); -3 doesn't equal -20 so the first equation isn't the right one. do you understand?